# Preservation of inequalities under Hadamard products

### v1

@@ 4. Symmetric decompositions
Recall that for any polynomial $h$ of degree at most $d$ there exist unique symmetric polynomials $a$ and $b$ such that $h(x)=a(x)+x\,b(x)$, $a=\mathcal{I}_{d}(a)$, and $b(x)=\mathcal{I}_{d-1}(b)$. The pair $(a,b)$ is called the symmetric decomposition or $\mathcal{I}_{d}$-decomposition of $h$
@@
[BS21, Lemma 2.2] Let $f$ be a polynomial of degree at most $d$. Then there exist unique polynomials $\widetilde{a}$ and $\widetilde{b}$ of degree at most $d$ and $d-1$, respectively, such that $\mathscr{R}_{d}(\widetilde{a})=\widetilde{a}$, $\mathscr{R}(\widetilde{b})=\widetilde{b}$ and
$f\ =\ \widetilde{a}+x\widetilde{b}\,.$
@@
The pair $(\widetilde{a},\widetilde{b})$ in the previous lemma is called the (symmetric) $\mathscr{R}_{d}$-decomposition of the polynomial $f$. The next lemma clarifies the relation between the $\mathcal{I}_{d}$-decomposition and the $\mathcal{R}_{d}$-decomposition.
@@
[BS21, Lemma 2.3] Let $h$ be a polynomial of degree at most $d$ and let $f=f(h;x)$ be its $f$-polynomial (with respect to $d$). Let $(a,b)$ be the $\mathcal{I}_{d}$-decomposition of $h$ and $(\widetilde{a},\widetilde{b})$ be the $\mathscr{R}_{d}$-decomposition of $f$. Then
@@
That is, $\widetilde{a}=f(a;x)$ and $\widetilde{b}=f(b;x)$ are the $f$-polynomials of $a$ and $b$ (with respect to $d$, respectively, $d-1$).
@@ 4.1. Nonnegative symmetric decompositions
The symmetric decomposition $(a,b)$ of $h=\mathcal{W}(p)$ is called nonnegative if both $a$ and $b$ have only nonnegative coefficients. An immediate consequence of Lemma 4.2 is that $h=\mathscr{W}(p)$ has a nonnegative symmetric decomposition if and only if the $f$-polynomial $f(h;x)=\mathscr{E}(p)$ has an magic positive $\mathscr{R}_{d}$-decomposition $(\widetilde{a},\widetilde{b})$, i.e., if both $\widetilde{a}$ and $\widetilde{b}$ are magic positive. As the following lemma shows, the property of magic positivity is preserved under basic operations. Observe that we recover the preservation of nonnegative coefficients under the Hadamard product which was previously shown in [FK14, Proposition 5.1 (i)].
@@
Since $\mathscr{R}_{d_{i}-1}(\widetilde{b}_{i})=\widetilde{b}_{i}$, it follows that $h_{j}^{i}=h_{d_{i}-1-j}^{i}$ for all $10\leq j\leq d_{i}-$ 1. That is, $\mathscr{W}(p_{i})$ is symmetric with defect $\operatorname{def}\mathscr{W}(p_{i})=1$ for $i=1,2$. By Theorem 1.3, also $\mathscr{W}(p_{1}p_{2})$ is symmetric with defect $\operatorname{def}\mathscr{W}(p_{1}p_{2})=1$. Thus, the corresponding $f$-polynomial is of the form
@@
with $c_{j}=c_{d_{1}+d_{2}-1-j}$ for $10\leq j\leq d_{1}+d_{2}-$ 1. Hence the first claim follows by setting $\ell=\sum_{j=0}^{d_{1}+d_{2}-1}c_{j}x^{j}(x+1)^{d_{1}+d_{2}-1-j}$. For the second claim we observe that
@@
We will show equivalently that if the $f$-polynomials $f_{1},f_{2}$ have magic positive $\mathscr{R}_{d_{1}}$- and $\mathscr{R}_{d_{2}}$-decompositions, then $f_{1}\diamond f_{2}$ has a magic positive $\mathscr{R}_{d_{1}+d_{2}}$-decomposition. Let $(\widetilde{a}_{i},\widetilde{b}_{i})$ be the decomposition of $f_{i}$, $i=1,2$, and let $(\widetilde{c},\widetilde{d})$ be the decomposition of $f_{1}\diamond f_{2}$. Then by Lemma 4.1
@@
$\widetilde{c}\enspace$ $=\enspace\widetilde{f}_{1}\diamond\widetilde{f}_{2}-x\widetilde{d}$
$=\enspace\widetilde{a}_{1}\diamond\widetilde{a}_{2}+(x\widetilde{b}_{1})\diamond\widetilde{a}_{2}+\widetilde{a}_{1}\diamond(x\widetilde{b}_{2})+(x\widetilde{b}_{1})\diamond(x\widetilde{b}_{2})$
@@
$=\enspace\widetilde{a}_{1}\diamond\widetilde{a}_{2}+\left[\widetilde{a}_{1}\diamond(x\widetilde{b}_{2})-x(\widetilde{a}_{1}\diamond\widetilde{b}_{2})\right]+\left[(x\widetilde{b}_{1})\diamond\widetilde{a}_{2}-x(\widetilde{b}_{1}\diamond\widetilde{a}_{2})\right]$
@@ 4.2. Interlacing symmetric decompositions
Following [BS21] we call the symmetric decomposition $(a,b)$ of a polynomial $h=\mathscr{W}(p)$ nonnegative if both $a$ and $b$ have only nonnegative coefficients, and interlacing if $b\preceq a$. Instead of working with the numerator polyomials directly, we will again consider the corresponding $f$-polynomial, $f(h;x)=\mathscr{E}(p)$. Crucially, by equation (4.1), $\mathscr{W}(p)$ and $\mathscr{E}(p)$ are related by a Möbius transformation that preserves real-rootedness of polynomials. In particular, $h$ has zeros in $(-\infty,0]$, i.e. only nonnegative coefficients, if and only if $f$ has zeros in $(-1,0]$. Moreover, the transformation preserves interlacing polynomials with zeros in these intervals. It follows that if $(\widetilde{a},\widetilde{b})$ is the $\mathscr{R}$-decomposition of $f(h;x)$ and the $\mathcal{I}_{d}$ composition $(a,b)$ of $h$ has only nonnegative coefficients then
@@
[Brä04, Theorem 12] Let $f,g$ and $h$ be real-rooted polynomials, and let $h$ have zeros in $[-1,0]$. Then $f\diamond h$ is real-rooted, and if $f\preceq g$ then
@@
We prove the equivalent statement for $f$-polynomials. We may thus assume that we have $f$-polynomials $f_{1},f_{2}$ of degrees $d_{1}$ and $d_{2}$ with magic positive $\mathscr{R}_{d_{1}}$- and $\mathscr{R}_{d_{2}}$-decompositions $(\widetilde{a}_{i},\widetilde{b}_{i})$ , $i=1,2$, and $\widetilde{b}_{i}\preceq\widetilde{a}_{i}$. Let $(\widetilde{c},\widetilde{d})$ be the $\mathscr{R}_{d_{1}+d_{2}}$-decomposition of $f_{1}\diamond f_{2}$. We will show that $\mathscr{R}(f_{1}\diamond f_{2})\preceq f_{1}\diamond f_{2}$.
Since $\widetilde{b}_{i}\preceq\widetilde{a}_{i}$, $\deg\widetilde{b}_{i}=\deg\widetilde{a}_{i}-1$, and the zeros of $\widetilde{b}_{i},\widetilde{a}_{i}$ lie in $[-1,0]$, for $i=1,2$, by Theorem 4.8 we have the following chain of interlacing polynomials.
@@ 5. Hadamard powers and their behavior in the limit
Clearly, for all integers $n\geq 0$, $|nP^{k}\cap\mathbb{Z}^{dk}|=|nP\cap\mathbb{Z}^{d}|^{k}$. Thus, we have $E_{P^{k}}(x)=E_{P}(x)^{k}$, and consequentially, by equation (4.3),
@@
We consider the $3$-dimensional Reeve tetrahedron with vertices $(0,0,0)$, $(1,0,0)$, $(0,1,0)$ and $(1,7,8)$ and show that the $f$-polynomial of the Ehrhart polynomial of any product with itself fails to be log-concave in the following theorem, thereby refuting Conjecture 1.6.
@@
$\enspace+(3+20x+24x^{2})\left(\sum_{\ell\geq 0}(\ell+1)f_{k,\ell+1}\,x^{\ell}\right)\,x\,(x+1)$
$\enspace+\frac{1}{4}(20+48x)\left(\sum_{\ell\geq 0}(\ell+2)(\ell+1)f_{k,\ell+2}\,x^{\ell}\right)\,x^{2}\,(x+1)^{2}$
$\enspace+\frac{1}{36}\cdot 48\cdot\left(\sum_{\ell\geq 0}(\ell+3)(\ell+2)(\ell+1)f_{k,\ell+3}\,x^{\ell}\right)\,x^{3}\,(x+1)^{3}.$
@@
$f_{k,0}=1,\qquad f_{k,1}=4^{k}-1,\qquad\text{ and }\enspace f_{k,2}=17^{k}-2\cdot 4^{k}+1\,.$
@@ 6. Final remarks
This question was considered before by Ferroni and Higashitani [FH24, Question 3.4(i)] in the context of Ehrhart polynomials. By an example provided in that article [FH24, Example 3.5], it is known that if $\mathscr{W}(p)$ and $\mathscr{W}(q)$ are nonnegative and unimodal, it is not necessarily true that $\mathscr{W}(pq)$ is unimodal too. Recently, Balletti [Bal23] showed that counterexamples coming from Ehrhart polynomials of lattice polytopes also exist in very large dimensions.
@@
[…]
@@
Notice, however, that a necessary step in the proof of Theorem 1.2 was that if $p(x,y)$ is Lorentzian, then $p(x+z,y+z)$ is Lorentzian. An analogous statement does not hold for denormalized Lorentzian polynomials. For example, the polynomial $p(x,y)=x^{2}+5xy+25y^{2}$ is denormalized Lorentzian, but $p(x+z,y+z)$ is not denormalized Lorentzian. In particular, there is no straightforward modification of our argument yielding a positive answer to Question 6.1. On the other hand, it would also be of interest to prove that the operator defined in Lemma 3.2 preserves the denormalized Lorentzian property. A proof of that fact would imply that if $\gamma_{f}$ is log-concave with no internal zeros, then $f$ is log-concave with no internal zeros. We have not been able to find counterexamples to that implication.
Acknowledgements: PB is a Wallenberg Academy Fellow supported by the Knut and Alice Wallenberg Foundation. LF was partially supported by grant 2018-03968 of the Swedish Research Council. KJ was partially supported by the Wallenberg AI, Autonomous Systems and Software Program funded by the Knut and Alice Wallenberg Foundation as well as grant 2018-03968 of the Swedish Research Council.

### v2

@@ 4. Symmetric decompositions
Recall that for any polynomial $h$ of degree at most $d$ there exist unique symmetric polynomials $a$ and $b$ such that h(x)=a(x)+xb(x)h(x)=a(x)+x\,b(x), $a=\mathcal{I}_{d}(a)$, and $b(x)=\mathcal{I}_{d-1}(b)$. The pair $(a,b)$ is called the symmetric decomposition or $\mathcal{I}_{d}$-decomposition of $h$ .
@@
[BS21, Lemma 2.2] Let $f$ be a polynomial of degree at most $d$. Then there exist unique polynomials $\widetilde{a}$ and $\widetilde{b}$ of degree at most $d$ and d1d-1, respectively, such that $\mathscr{R}_{d}(\widetilde{a})=\widetilde{a}$, $\mathscr{R}(\widetilde{b})=\widetilde{b}$ and
[…]
@@
[…]
@@
[…]
@@
That is, $\widetilde{a}=f(a;x)$ and $\widetilde{b}=f(b;x)$ are the $f$-polynomials of $a$ and $b$ (with respect to $d$, respectively, d1d-1).
@@ 4.1. Nonnegative symmetric decompositions
$\text{$\mathscr{W}(p)$ has nonnegative coefficients}\iff\text{$\mathscr{E}(p)$ is magic positive}$
@@
Since $\mathscr{R}_{d_{i}-1}(\widetilde{b}_{i})=\widetilde{b}_{i}$, it follows that $h_{j}^{i}=h_{d_{i}-1-j}^{i}$ for all $0\leq j\leq d_{i}-1$. That is, $\mathscr{W}(p_{i})$ is symmetric with defect $\operatorname{def}\mathscr{W}(p_{i})=1$ for $i=1,2$. By Theorem 1.3, also $\mathscr{W}(p_{1}p_{2})$ is symmetric with defect $\operatorname{def}\mathscr{W}(p_{1}p_{2})=1$. Thus, the corresponding $f$-polynomial is of the form
@@
with $c_{j}=c_{d_{1}+d_{2}-1-j}$ for $0\leq j\leq d_{1}+d_{2}-1$. Hence the first claim follows by setting $\ell=\sum_{j=0}^{d_{1}+d_{2}-1}c_{j}x^{j}(x+1)^{d_{1}+d_{2}-1-j}$. For the second claim we observe that
@@
[…]
@@
$\widetilde{c}\kern 5.0pt$ $=\kern 5.0pt\widetilde{f}_{1}\diamond\widetilde{f}_{2}-x\widetilde{d}$
$=\kern 5.0pt\widetilde{a}_{1}\diamond\widetilde{a}_{2}+(x\widetilde{b}_{1})\diamond\widetilde{a}_{2}+\widetilde{a}_{1}\diamond(x\widetilde{b}_{2})+(x\widetilde{b}_{1})\diamond(x\widetilde{b}_{2})$
@@
$=\kern 5.0pt\widetilde{a}_{1}\diamond\widetilde{a}_{2}+\left[\widetilde{a}_{1}\diamond(x\widetilde{b}_{2})-x(\widetilde{a}_{1}\diamond\widetilde{b}_{2})\right]+\left[(x\widetilde{b}_{1})\diamond\widetilde{a}_{2}-x(\widetilde{b}_{1}\diamond\widetilde{a}_{2})\right]$
@@ 4.2. Interlacing symmetric decompositions
[…]
@@
[Brä04, Theorem 12] Let $f,g$ and $h$ be real-rooted polynomials, and let $h$ have zeros in [1,0][-1,0]. Then $f\diamond h$ is real-rooted, and if $f\preceq g$ then
@@
[…]
Since $\widetilde{b}_{i}\preceq\widetilde{a}_{i}$, $\deg\widetilde{b}_{i}=\deg\widetilde{a}_{i}-1$, and the zeros of $\widetilde{b}_{i},\widetilde{a}_{i}$ lie in [1,0][-1,0], for $i=1,2$, by Theorem 4.8 we have the following chain of interlacing polynomials.
@@ 5. Hadamard powers and their behavior in the limit
Clearly, for all integers $n\geq 0$, one has $|nP^{k}\cap\mathbb{Z}^{dk}|=|nP\cap\mathbb{Z}^{d}|^{k}$. Therefore, $E_{P^{k}}(x)=E_{P}(x)^{k}$ and, consequentially, by equation (4.3),
@@
We consider the $3$-dimensional Reeve tetrahedron with vertices $(0,0,0)$, $(1,0,0)$, $(0,1,0)$ and $(1,7,8)$ and show that the $f$-polynomial of the Ehrhart polynomial of all powers of it fail to be log-concave, hence refuting Conjecture 1.6.
@@
$\kern 5.0pt+(3+20x+24x^{2})\left(\sum_{\ell\geq 0}(\ell+1)f_{k,\ell+1}\,x^{\ell}\right)\,x\,(x+1)$
$\kern 5.0pt+\frac{1}{4}(20+48x)\left(\sum_{\ell\geq 0}(\ell+2)(\ell+1)f_{k,\ell+2}\,x^{\ell}\right)\,x^{2}\,(x+1)^{2}$
$\kern 5.0pt+\frac{1}{36}\cdot 48\cdot\left(\sum_{\ell\geq 0}(\ell+3)(\ell+2)(\ell+1)f_{k,\ell+3}\,x^{\ell}\right)\,x^{3}\,(x+1)^{3}.$
@@
$f_{k,0}=1,\qquad f_{k,1}=4^{k}-1,\qquad\text{ and }\kern 5.0ptf_{k,2}=17^{k}-2\cdot 4^{k}+1\,.$
@@ 6. Final remarks
This question was considered before by Ferroni and Higashitani [FH24, Question 3.4(i)] in the context of Ehrhart polynomials. By an example provided in that article [FH24, Example 3.5], it is known that if $\mathscr{W}(p)$ and $\mathscr{W}(q)$ are nonnegative and unimodal, it is not necessarily true that $\mathscr{W}(pq)$ is unimodal too. Recently, Balletti [Bal24] showed that counterexamples coming from Ehrhart polynomials of lattice polytopes also exist in very large dimensions.
@@
$p\in\mathrm{H}_{2}^{n}\text{ is denormalized Lorentzian}\iff$ $p(x,1)\text{ is log-concave without internal zeros.}$
@@
Notice, however, that a necessary step in the proof of Theorem 1.2 was that if $p(x,y)$ is Lorentzian, then $p(x+z,y+z)$ is Lorentzian. An analogous statement does not hold for denormalized Lorentzian polynomials. For example, the polynomial $p(x,y)=x^{2}+5xy+25y^{2}$ is denormalized Lorentzian, but $p(x+z,y+z)$ is not denormalized Lorentzian. In particular, there is no straightforward modification of our argument yielding a positive answer to Question 6.1. On the other hand, it would also be of interest to prove that the operator defined in Lemma 3.3 preserves the denormalized Lorentzian property. A proof of that fact would imply that if $\gamma_{f}$ is log-concave with no internal zeros, then $f$ is log-concave with no internal zeros. [^1] After the posting of the present paper on the arXiv, Ferroni, Panova, and Venturello [FPV25] proved this implication.
Acknowledgements: PB is a Wallenberg Academy Scholar supported by the Knut and Alice Wallenberg Foundation, and the Göran Gustafsson foundation. LF was partially supported by grant 2018-03968 of the Swedish Research Council, and by the Minerva Research Foundation at the Institute for Advanced Study. KJ was partially supported by the Wallenberg AI, Autonomous Systems and Software Program funded by the Knut and Alice Wallenberg Foundation as well as grants 2018-03968 and 2023-04063 of the Swedish Research Council, and the Verg Foundation.